Anna is taking a test in which items of type A are worth 10 points and items of type B are worth 15 points. It takes 3 minutes to complete each item of type A and 6 minutes to complete each item of type B. The total time allowed is 60 minutes and Anna answers exactly 16 questions. How many questions of each type did she complete? Assuming that all her answers were correct, what was her score?
Answer by macston(5194)

A+B=16 Multiply this equation by3
3A+3B=48
3A+6B=60 Subtract from above.
-3B=-12 Divide each side by -3.
B=4 ANSWER 1: She answered 4 questions of type B.
A+B=16
A+4=16
A=12 ANSWER 2: She answered 12 type A questions.
.
10A+15B=Score
10(12)+15(4)=Score
120+60=Score
180=Score ANSWER 3: Anna's score was 180.